Blue, Inc. is considering borrowing $100,000 at 5%. Its tax rate is 35%. What is the present value of the interest tax shield if the debt is permanent? (do not use the $ sign when entering your answer).
Answer #1

Dear student thankyou for using HomeworkLib. If you love the answer then please gives thumbsup.

Calculation of present value of interest tax shield:

Tax rate= 35%

Value of borrowing= 100000

Interest= 100000*0.05= 5000

Annual tax shield= Interest * tax rate= 5000*0.35= 1750

Present value of interest tax shield= annual tax shield/interest rate

Present value of interest tax shield= 1750/0.05= 35000

Present value of interest tax shield is 35000